OpenStudy (anonymous):

A local market wants to mix some $5 apples with some $4 oranges. (These cost are per pountd) How many pounds of each item should they mix to obtain a 30 pound mixture that sells for $4.40 per pound?

OpenStudy (anonymous):

do i use elimination?

OpenStudy (whpalmer4):

you can solve the equations with substitution or elimination...the trick is setting up the equations in the first place.

OpenStudy (jhannybean):

I had written it wrong earlier, sorry about that.

OpenStudy (anonymous):

wait, so is it x+y=30 5x+4y=4.40?

OpenStudy (whpalmer4):

no, it has to be 5x + 4y = 30*4.40 — the value of the fruit is what the total amount will sell for.

OpenStudy (anonymous):

sorry but i don't quite get it .....

OpenStudy (whpalmer4):

they sell apples for $5/lb. they sell oranges for $4/lb. they want to put together 30 lbs of fruit in a mixture that they can sell for $4.40/lb. The total selling price is 30*$4.40, just as if they had sold them separately in the proper proportions. so, if x is the number of apples, and y is the number of oranges, $5/lb *x + $4/lb * y = 30*$4.40.

OpenStudy (anonymous):

so oranges=18 and apples=12?

OpenStudy (whpalmer4):

let's check: 18+12 = 30 so far so good $5/lb * 12 lb + $4/lb *18 lb = 30 lb*$4.40/lb $60 + $72 = $132 looks good!

OpenStudy (anonymous):

all right. thanks for the help :)